# The exact travelling wave solutions of a KPP equation

### v3

@@ III The ODE which doesn’t contain explicitly the independent variable
where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
$u_{\xi\xi}+cu_{\xi}$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (12a)
$pp_{u}+cp$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (12b)
Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (12a) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath . One can easily check up that there exists the solution of (12b) satisfying the boundary conditions (10)
@@
Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12a) as
@@
In Appendix A we’ll see how the results (13) - (16) can be generalized in the case when advection is taken into account in addition to diffusion. In Appendix C we’ll analyze the solution of (12b) (for $n=1$) for $c$ close to $c_{0}$.
@@
In the previous Section we postulated two boundary conditions (10) for (9). However, the equation is of the first order. The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s. of the equation. So one may think that we should finetune the speed $c$ for the solution of (12b) to exist, and the solution will not exist for a general value of $c$. In fact, didn’t we finetune $c$ in the previous Section?
However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (12b) for arbitrary $c\gg 1$ is presented. It looks like we have a problem. We have another problem with the boundary conditions (10) - they are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$. In this Section we’ll solve both problems. We’ll show that one of the boundary conditions (10) should be specified, and the other can be discarded.
Consider the solution of Eq. (5) in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
$u_{\xi}$ $=p,$ (18a)
$p_{\xi}$ $=-cp-F^{\prime}(0)u.$ (18b)
In the second case, the equation can be presented as
$v_{\xi}$ $=-p,$ (19a)
$p_{\xi}$ $=cp-|F^{\prime}(1)|v,$ (19b)
@@
Let us start from analysing (19). The system can be reduced to the canonical form
$X_{\xi}$ $=s_{1}X,$ (20a)
$Y_{\xi}$ $=s_{2}Y,$ (20b)
where
$X$ $=p+s_{2}v,$ (21a)
$Y$ $=p+s_{1}v,$ (21b)
and $s_{1,2}$ are the roots of the characteristic equation
@@
$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (23)
(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.)
Because $u(\xi)$ should satisfy the boundary condition (6a), we obtain
$Y=0.$ (24)
Returning to Eq. (9) we realize that the boundary condition (10a) should be specified as follows: In the vicinity of $u=1$
@@
Now let us come to (18). The characteristic equation for the system is
$s^{2}+cs+F^{\prime}(0)=0.$ (26)
The roots of the characteristic equations are
$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4F^{\prime}(0)}}{2}.$ (27)
We have to consider separately the case of the real roots, the case of the complex roots and the case of the double root.
Let us start from the case of the real roots ($c^{2}>4F^{\prime}(0)$). In this case we can reduce the system (18) to
$\frac{dX}{s_{3}X}=\frac{dY}{s_{4}Y},$ (28)
where
$X$ $=p+s_{4}v,$ (29a)
$Y$ $=p+s_{3}v.$ (29b)
The general integral is given by the equation
$Y=CX^{\frac{s_{4}}{s_{3}}}.$ (30)
Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node. (Note that $c$ given by (15) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
For $c^{2}<4F^{\prime}(0)$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (18) to the form goursat
$\frac{dX}{\alpha X-\beta Y}=\frac{dY}{\beta X+\alpha Y}$ (31)
from which we derive
$\frac{XdX+YdY}{\alpha\left(X^{2}+Y^{2}\right)}=\frac{XdY-YdX}{\beta\left(X^{2}+Y^{2}\right)}$ (32)
The general integral of the equation (18) is therefore represented by the equation
$\sqrt{X^{2}+Y^{2}}=Ce^{\frac{\alpha}{\beta}\tan^{-1}\frac{Y}{X}}.$ (33)
All these curves have the form of spirals which approach the origin as an asymptotic point. The origin is said to he a focus.
If the characteristic equation has a double root $s=c/2$, which occurs for $c^{2}=4F^{\prime}(0)$, the system (18) reduces to the form
$\frac{dX}{X}=\frac{dY}{X+Y},$ (34)
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e^{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus.
Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable. (Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated (saarloos ; ma .)
Now we understand that the second boundary condition (10b) can be discarded. It is satisfied automatically. And returning to the previous Section: we did finetune the speed $c$ for the solution of (12b) to be expressed in terms of elementary functions, and not for the solution to exist.
@@ Acknowledgements.
We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed and C. Ou for the insightful discussions.
@@ Appendix A Generalized Burgers–Huxley equation
The results obtained in Section II for the generalized KPP equation (12a) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
@@
For the travelling wave solution of (35) we obtain instead of (12b) the equation
@@ Appendix C Perturbation theory II
In Section III we have solved Eq. (12b) for $c=c_{0}$. Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution as

### v4

@@ III The ODE which doesn’t contain explicitly the independent variable
where $n>0$ and $a\geq 0$ are some constants, so (5) takes the form
$u_{\xi\xi}+cu_{\xi}=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (12)
Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (12) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
Starting from (12), we obtain Eq. (9) in the form
$pp_{u}+cp=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (13)
One can easily check up that there exists the solution of (13) satisfying the boundary conditions (10)
@@
Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12) as
@@
In Appendix A we’ll see how the results (14) - (17) can be generalized in the case when advection is taken into account in addition to diffusion. In Appendix C we’ll analyze the solution of (13) (for $n=1$) for $c$ close to $c_{0}$.
@@
In the previous Section we postulated two boundary conditions (10) for (9). However, the equation is of the first order. The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s. of the equation. So one may think that we should finetune the speed $c$ for the solution of (13) to exist, and the solution will not exist for a general value of $c$. In fact, didn’t we finetune $c$ in the previous Section? However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (13) for arbitrary $c\gg 1$ is presented. It looks like we have a problem.
To solve this problem we have to return to Eq. (5) with the boundary conditions (6). Actually, the mechanical analogy allows us easily visualize the solution of the equation, but let us forget for the time being about this analogy and look at the equation purely mathematically. We immediately realize that the boundary conditions (6) - they are exceptional goursat $u_{\xi\xi}$ is not defined at $\xi=\pm\infty$. To make use of the boundary conditions consider Eq. (5) for $\xi\to\pm\infty$. For $\xi\to+\infty$ the equation can be presented as
$u_{\xi\xi}+cu_{\xi\xi}+F^{\prime}(0)u=0.$ (19)
For $\xi\to-\infty$ the equation can be presented as
$v_{\xi\xi}+cv_{\xi}-|F^{\prime}(1)|v=0,$ (20)
@@
The solution of (20) is
$v=Ae^{s_{1}\xi}+Be^{s_{2}\xi},$ (21)
where $A$ and $B$ are arbitrary constants, and $s_{1,2}$ are the roots of the characteristic equation
@@
$s_{1}$ $=\frac{c+\sqrt{c^{2}+4|F^{\prime}(1)|}}{2},$ (23a)
$s_{2}$ $=\frac{c-\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (23b)
Because of the boundary condition (6a), $B=0$. So this boundary condition should be specified to: For $\xi\to-\infty$
$1-u\sim e^{s_{1}\xi}.$ (24)
Now let us come to (19). The characteristic equation is
$s^{2}+cs+F^{\prime}(0)=0,$ (25)
that is
$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4F^{\prime}(0)}}{2}.$ (26)
We have either the real roots (both being negative), the double (negative) root, or the complex roots (both with the negative real part). In the first case solution of (19) is
$u=Ae^{s_{3}\xi}+Be^{s_{4}\xi},$ (27)
for the case of double root the solution is
$u=e^{-c\xi/2}(A+B\xi),$ (28)
and in the complex case the solution is
$u=e^{-c\xi/2}\left[A\cos(\alpha\xi)+B\sin(\alpha\xi)\right],$ (29)
where $\alpha=[4F^{\prime}(0)-c^{2}]/2$. In any case the solution of (19) satisfies the boundary condition (6b) for any $A$ and $B$. So the boundary condition (6b) can be discarded.
Because we have lost one boundary condition the solution of (5) contains an arbitrary constant. And here our mechanical analogy helps us to understand that this (irrelevant) constant describes the translation of the solution in ”time” $\xi$.
After we understood the true nature of the boundary conditions (6) for (5), the situation with Eq. (9) becomes obvious. Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
However, looking back at (24) we realize that the boundary condition (10a) should be specified as follows: For $u\to 1$
@@
And looking back at our analysis of (19) we realize that any solution of Eq. (9) will satisfy (10b), so the second boundary condition (10b) can be discarded. It is satisfied automatically.
So finally, in the previous Section we were solving the first order ODE with the single boundary condition. The solution should exist for any given value of the speed $c$ and shouldn’t contain an undefined constant. We did finetune the speed $c$ in the previous Section for the solution of (13) to be expressed in terms of elementary functions, and not for the solution to exist.
Our analysis of (19) allows us to understand (partially) the analytic properties of $p(u)$. If $c^{2}\geq 4F^{\prime}(0)$, from (27) and (28) we can expect that $p(u)$ would be either a single - or a double-valued function. (Further analysis is necessary to decide which option is realised.) Note that $c$ given by (16) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
Looking at (29) (which corresponds to $c^{2}<4F^{\prime}(0)$) on the other hand, we realize that $p(u)$ would be an infinitely-valued function. Note that if for physical reasons $u$ should be strictly nonnegative, such values of speed are unacceptable, because from (29) follows that $u$ would acquire negative values in the vicinity of $\xi=+\infty$. (Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated saarloos ; ma .)
@@ Acknowledgements.
We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed, C. Ou and R. Selvaraj for the insightful discussions.
@@ Appendix A Generalized Burgers–Huxley equation
The results obtained in Section II for the generalized KPP equation (12) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
@@
For the travelling wave solution of (31) we obtain instead of (13) the equation
@@ Appendix C Perturbation theory II
In Section III we have solved Eq. (13) for $c=c_{0}$. Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution as